# MonoMPC: Monocular Vision Based Navigation with Learned Collision Model and Risk-Aware Model Predictive Control

### v1

@@ Abstract
Navigating unknown environments with a single RGB camera is challenging, as the lack of depth information prevents reliable collision-checking. While some methods use estimated depth to build collision maps, we found that depth estimates from vision foundation models are too noisy for zero-shot navigation in cluttered environments.
We propose an alternative approach: instead of using noisy estimated depth for direct collision-checking, we use it as a rich context input to a learned collision model. This model predicts the distribution of minimum obstacle clearance that the robot can expect for a given control sequence. At inference, these predictions inform a risk-aware MPC planner that minimizes estimated collision risk. Our joint learning pipeline co-trains the collision model and risk metric using both safe and unsafe trajectories. Crucially, our joint-training ensures optimal variance in our collision model that improves navigation in highly cluttered environments. Consequently, real-world experiments show 9x and 7x improvements in success rates over NoMaD and the ROS stack, respectively. Ablation studies further validate the effectiveness of our design choices.
@@ I Introduction
Recent progress in vision foundation models like DepthAnything [2, 1] and ZoeDepth [3] has enabled depth prediction from a single RGB image, opening new avenues for monocular navigation. Prior work has used these depth estimates to build local 3D or collision maps for planning [4]. However, our experiments show that such depth predictions are often too noisy or inconsistent for reliable collision-checking, particularly in close-proximity scenarios where small errors can result in collisions (see Fig. LABEL:fig:teaser). In zero-shot settings—where the robot must generalize to unseen environments—these models fall short as standalone solutions.
In this work, we propose a novel approach to monocular navigation that reinterprets estimated depth not as a proxy for ground truth, but as an informative input to a learned probabilistic collision model. While monocular depth estimates may be noisy, they contain valuable structural cues. Our model takes the RGB image and a candidate control sequence as input and predicts a distribution over worst-case obstacle clearance along that trajectory. These predictions are used at inference time to estimate collision risk, which is minimized via a risk-aware model predictive control (MPC) planner to generate safe and goal-directed actions.
We propose a joint learning pipeline that co-trains the collision model and risk metric using both safe and unsafe trajectories. This ensures the model learns how its predictions impact downstream risk, providing valuable feedback that helps regularize variance to enable safe yet non-conservative behavior.
We validate our approach through real-world hardware experiments, showing substantial improvements in navigation success over strong baselines like NoMaD [5]—a diffusion-based end-to-end method—and the ROS navigation stack, which relies on estimated depth-based costmaps (Fig. LABEL:fig:teaser). Ablation studies further highlight the impact of our design choices.
@@ II-A End-to-End Approaches for Visual Navigation
A common approach in monocular visual navigation is to train end-to-end models that map RGB images directly to control outputs like velocity commands or waypoints. Many rely on supervised learning via behavioral cloning or imitation learning [6][7][8][9][10][11]. While simple and appealing, these methods often generalize poorly, are sensitive to perception noise, and lack interpretability. They also omit explicit modeling of geometry, uncertainty, and constraints, making them data-hungry and hard to adapt. Recent methods like ViNT [12] and NoMaD [5] improve generalization by using image-conditioned diffusion models to generate trajectories from RGB input. However, they depend on high-quality supervision and do not model collision uncertainty.
@@
Another class of approaches focuses on learning spatial representations from monocular input [9, 10, 4]. For example, [4] constructs a map from estimated depth and then performs planning over it. However, since our problem setting focuses on navigation in unknown environments without explicit map construction, such methods fall outside the scope of this work.
@@
Deep reinforcement learning (DRL) has also been explored for visual navigation, particularly in unseen or partially known environments [13][14][15]. While DRL enables flexible end-to-end learning from experience, it typically requires extensive training in simulation, careful reward shaping, and offers limited safety guarantees.
II-B World Model-Based Approaches for Visual Navigation
@@
World model-based methods [16][17][18] learn action-conditioned predictive models to simulate dynamics—predicting the next state given the current state and action—for downstream planning and control, rather than directly learning policies.  [19] adopt a slightly different approach and propose a visual MPC controller that outputs a sequence of velocity commands based on the current image and a trajectory of subgoal images. Authors in [20] regress depth images and actions to collision probabilities which is then minimized within an MPC framework. Although theoretically, this approach can be extended to work with estimated depth, it is unclear how the associated noise will affect the efficiency of the overall pipeline. Moreover, we believe that predicting obstacle clearance distribution is easier than directly predicting collision probabilities, as the supervision data for the former can be easily obtained. Our design choice also allows us to use sophisticated statistical tools to capture collision risk from the clearance samples in an efficient manner.
II-C Improvement Over State-of-the-Art:
@@
Unlike end-to-end pipelines, our approach explicitly incorporates safety constraints by learning an action-conditioned collision model, placing us closer to world-model methods. However, unlike [16][17][18], [19], we model predictive uncertainty and collision risk, and introduce a novel method to learn task-aware uncertainty. Moreover, we also learn optimal parameters of our risk metric to improve the efficacy of our approach.
Most monocular navigation methods demonstrate results in simple environments like hallways. In contrast, we show—for the first time—robust and reliable navigation in cluttered settings typically requiring LiDAR or depth sensors.
@@ III-A Vision-Based Navigation as Risk-Aware Trajectory Optimization
$\mathbf{x}_{k+1}=f(\mathbf{x}_{k},\mathbf{u}_{k}),$ (1b)
$\mathbf{u}_{min}\leq\mathbf{u}_{k}\leq\mathbf{u}_{max}\forall k$ (1c)
where $c(\cdot)$ represents the state-dependent cost function. The vector $\mathbf{x}_{k}$ represents the state of the robot at time-step $k$. The vector $\mathbf{x}$ is the concatenation of the states at different $k$. The function $f$ represents the robot dynamics model. The control inputs are represented by $\mathbf{u}_{k}$. The vector $\mathbf{u}$ is formed by stacking $\mathbf{u}_{k}$ at different time-step $k$ respectively. The scalar function $d$ represents a learned probabilistic collision model, which takes as input the observations $\mathbf{o}$ and the control sequence $\mathbf{u}$ and outputs a distribution of worst-case obstacle clearance along the trajectory resulting from $\mathbf{u}$. We discuss the details of the $d$ in the subsequent sections. The first term in (1a) minimizes the state cost, typically addressing path-following errors. The second term captures collision-risk based on the predictions of $d$. The last term in (1a) penalizes large control inputs, and the weights $w_{i}$ tune the robot’s risk-seeking behavior. Control bounds are enforced through (1c). We construct a MPC feedback loop by solving (1a)-(1c) in a receding horizon manner from the current state.
@@ III-B Modeling Risk Through Chance-Constraints
$r(h(d(\mathbf{u},\mathbf{o})))=P(h(d(\mathbf{u},\mathbf{o})))\geq 0)$ (2)
Intuitively, our risk model captures the probability of collision constraints being violated for a given control sequence $\mathbf{u}$ and current observation $\mathbf{o}$. Moreover, the risk depends on the distribution characterized by the collision model $d$. Our aim is to first learn a suitable $d$ and then minimize the associated risk using (1a)-(1c).
A key challenge in using risk $r$ as defined in (2) is that the left-hand side does not have an analytical form if the predictive distribution of $d(\mathbf{u},\mathbf{o}))$ departs significantly from Gaussian. Thus, in the next subsection, we present a surrogate risk model that is more tractable and can produce the same effect as $r$.
@@ III-C Maximum Mean Discrepancy(MMD) as parameterized risk metric
In the deterministic scenario, driving $\overline{h}(d(\mathbf{u},\mathbf{o}))$ to zero will push $h(d(\mathbf{u},\mathbf{o}))$ to the feasible boundary. In the stochastic case, (3) maps $d(\mathbf{u},\mathbf{o})$ to a distribution of constraint residuals. Let $\overline{h}(d(\mathbf{u},\mathbf{o}))\sim p_{\overline{h}}$. Although we don’t know the parametric form for $p_{\overline{h}}$, we can be certain that its entire mass lies to the right of $\overline{h}=0$. Moreover, as $P(h(d(\mathbf{u},\mathbf{o}))\geq 0)$ approaches zero, $p_{\overline{h}}$ converges to a Dirac-Delta distribution $p_{\delta}$ [21]. In other words, one way of reducing risk is to minimize the difference between $p_{\overline{h}}$ and $p_{\delta}$. We formulate this distribution matching through the lens of Maximum Mean Discrepancy (MMD) [22],[21] and use the following surrogate for the collision-risk proposed in [21], [23].
@@
The $K_{\lambda}$ is a positive definite function in RKHS known as the kernel function with the ”reproducing” property (6) [22] , where $\left\langle\cdot\right\rangle$ represents the inner product. Throughout this paper, we have used Laplacian kernel for which $\lambda$ represents the kernel width.
@@ III-D Learning the Collision Model: The Baseline Approach
This section presents our approach for learning a probabilistic collision model $d(\mathbf{u},\mathbf{o})$ which quantifies the worst-case obstacle clearance along a given trajectory. We model it as a Gaussian random variable with mean $\mu_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$ and variance $\sigma_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$, both of which are designed as feed-forward neural networks parameterized by weight $\boldsymbol{\theta}$. Our design choice allows us to capture the noise in the observation $\mathbf{o}$ primarily stemming from the estimated point-cloud.
The detailed architecture is shown in Fig. 1. It consists of a pre-trained depth estimator that takes in RGB image and produces a depth-map. The depth-map is converted to a 2D point cloud and passed through PointNet++ [24] to get a feature embedding. We concatenate PointNet++ features with the initial state to get the full observation vector $\mathbf{o}$, which is then passed along with control sequence $\mathbf{u}$ to a Multi-Layer Perceptron (MLP) to get $\mu_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$ and $\sigma_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$. Both PointNet++ and MLP are trained in end-to-end manner using the following negative log-likelihood (NLL) function.
@@
1 $M$ = Maximum number of iterations
@@ 24 $cost\leftarrow$ $c_{q}$, $\forall q$ in the ConstraintEliteSetConstraintEliteSet
29 $({{}^{m+1}}\boldsymbol{\nu},{{}^{m+1}}\boldsymbol{\Sigma})\leftarrow$ Update distribution based on EliteSetEliteSet
@@
Algorithm 1 Sampling-Based Optimizer to Solve (1a)-(1c)
@@ III-E Learning the Collision Model:The Down-Stream Task-Aware Approach
The risk estimation head estimates collision constraint violations from $^{i}{d}$ ((8b)-(8c)). Using these violations and the predicted kernel width $\lambda_{\boldsymbol{\theta}}$, the empirical MMD risk $r_{MMD}^{emp}$ is computed. This risk is then passed through an MLP to yield a 2D feature vector $\mathbf{z}$, which a softmax layer normalizes to $\hat{\mathbf{y}}$. The ground-truth label for $\hat{\mathbf{y}}$ is given by $\mathbf{y}$ and is computed through (9)
@@
$^{i}h(^{i}d)=-{{}^{i}}{d}+d_{o}$ (8b)
$^{i}\overline{h}(^{i}d)=\max(0,{{}^{i}}h(^{i}d))$ (8c)
$\mathbf{y}=\begin{cases}[0,1]^{T},&\text{if }-d_{gt}+d_{o}\leq 0\\ [1,0]^{T},&\text{if }-d_{gt}+d_{o}>0\end{cases}$ (9)
The architecture of Fig.2 is trained with a combination of NLL and a cross-entropy loss (10). The latter acts as an implicit task-aware regularization on the predicted $\sigma_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$.
@@ III-F Overall Approach
Alg.1 presents our overall approach wherein we solve (1a)-(1c) using the trained neural worst-case obstacle clearance model as our collision predictor. We adopt a sampling-based optimizer based on [25], wherein a batch of control sequences are drawn from a distribution which is gradually refined across iterations.
The algorithm initializes the sampling distribution (Line 2) and samples control inputs (Line 5), which are passed through a trained neural network (Line 6) to predict the mean $\mu_{\boldsymbol{\theta},q}$ and variance $\sigma_{\boldsymbol{\theta},q}$ associated with the distribution of obstacle clearances, as well as the kernel hyperparameter $\lambda_{\boldsymbol{\theta},q}$. These predictions generate $N$ distance samples $^{i}d_{q}$, constraint violations $^{i}\overline{h}_{q}$ (Line 7), and subsequently $r_{MMD,q}^{emp}$ (Line 8). State trajectories $\mathbf{x}_{q}(\mathbf{u}_{q})$ are then generated using the robot dynamics (Line 9). The $n_{c}$ lowest-risk samples form the $ConstraintElliteSet$ (Line 10), from which costs are evaluated and stored (Lines 11–13). Finally, the top $n_{e}$ samples form the $ElliteSet$ (Line 14) used to update the sampling distribution (Line 15) following [25].
Figure 3: We evaluate our learned collision model’s and risk-aware MPC performance in a simple scenario. In (a), although the estimated point cloud (green) appears farther than the ground truth (red), the robot slows down (e) and steers away from the obstacle (a–d), successfully avoiding collision. This demonstrates our learned collision model’s ability to output a well-regularized distribution over worst-case obstacle clearance. Note that in (b) and (c) the obstacle is outside the field-of-view of the robot and hence the resulting point clouds are absent.

### v2

@@ Abstract
Navigating unknown environments with a single RGB camera is challenging, as the lack of depth information prevents reliable collision-checking. While some methods use estimated depth to build collision maps, we found that depth estimates from vision foundation models are too noisy for zero-shot navigation in cluttered environments. We propose an alternative approach: instead of using noisy estimated depth for direct collision-checking, we use it as a rich context input to a learned collision model. This model predicts the distribution of minimum obstacle clearance that the robot can expect for a given control sequence. At inference, these predictions inform a risk-aware MPC planner that minimizes estimated collision risk. We proposed a joint learning pipeline that co-trains the collision model and risk metric using both safe and unsafe trajectories. Crucially, our joint-training ensures well calibrated uncertainty in our collision model that improves navigation in highly cluttered environments. Consequently, real-world experiments show reductions in collision-rate and improvements in goal reaching and speed over several strong baselines.
@@ I Introduction
Recent progress in vision foundation models like DepthAnything [depthanything, NEURIPS2024_26cfdcd8] and ZoeDepth [https://doi.org/10.48550/arxiv.2302.12288] has enabled depth prediction from a single RGB image, opening new avenues for monocular navigation. Prior work has used these depth estimates to build local 3D or collision maps for planning [simon2023mononav]. However, depth estimates from existing models [depthanything, NEURIPS2024_26cfdcd8], [https://doi.org/10.48550/arxiv.2302.12288] are often too noisy or inconsistent for reliable collision-checking, particularly in close-proximity scenarios where small errors can result in collisions (see Fig. LABEL:fig:teaser). In zero-shot settings, where the robot must generalize to unseen environments, these models fall short as standalone solutions and we further discuss these observation in Section IV.
Contributions: This work introduces a principled framework for autonomous navigation that, for the first time, directly reasons about uncertainty from estimated depth. At the core of our approach is a learned collision model that predicts a distribution over the minimum obstacle clearance for a planned trajectory. By conditioning this model on depth estimates, we can quantify collision risk at inference time. This risk is then integrated into a risk-aware Model Predictive Control (MPC) planner, which optimizes for actions that are both safe and goal-oriented.
A core innovation of our approach lies in appropriately transferring the noise in depth estimation to the actual downstream collision risk estimate. This is achieved by developing a joint learning pipeline that co-trains the collision model and the risk metric using both safe and un-safe trajectories. Our approach ensures that during the training process, the probabilistic collision model is aware of how its predictions impact downstream risk. This in turn, provides crucial supervision signal during training that regularizes the variance of the collision model and enables safe yet non-conservative behavior.
We validate our approach on real-world hardware, showing significant navigation gains over two baseline groups. The first, including the ROS stack and MonoNav [simon2023mononav], builds occupancy maps from noisy depth estimates, often causing unsafe or overly conservative behavior (Fig. LABEL:fig:teaser). The second, NoMaD [10610665]-a diffusion-based end-to-end policy that directly regresses robot trajectory based on the input RGB image-lacks the precision needed for cluttered scenes. We also present additional statistical results validating the accuracy and statistical consistency of our learned collision model.
@@ II Related Works
End-to-End Approaches for Visual Navigation: A common approach in monocular visual navigation is to train end-to-end models that map RGB images directly to control outputs like velocity commands or waypoints. Many rely on supervised learning via behavioral cloning or imitation learning [8460487], [8264734], [chen2020learning], [chaplot2020learning], [chaplot2020neural], [Sadeghi-RSS-17]. While simple and appealing, these methods often generalize poorly, are sensitive to perception noise, and lack interpretability. They also omit explicit modeling of geometry, uncertainty, and constraints, making them data-hungry and hard to adapt. Recent methods like ViNT [shah2023vint] and NoMaD [10610665] improve generalization by using image-conditioned diffusion models to generate trajectories from RGB input. However, they depend on high-quality supervision and do not model collision uncertainty.
@@
Another class of approaches focuses on learning spatial representations from monocular input [chaplot2020learning, chaplot2020neural, simon2023mononav]. For example, [simon2023mononav] constructs a map from estimated depth on-the-fly and then performs planning over it. However, Section IV shows that errors in occupancy predictions due to noise in estimated depth proves detrimental in cluttered environments.
@@
Deep reinforcement learning (DRL) has also been explored for visual navigation, particularly in unseen or partially known environments [7989381], [8460655], [10629078]. While DRL enables flexible end-to-end learning from experience, it typically requires extensive training in simulation, careful reward shaping, and offers limited safety guarantees.
@@
Predictive Models for Visual Navigation: Approaches like [Bar_2025_CVPR], [hansen2024tdmpc2], [wu2022daydreamer] learn action and image conditioned predictive models to simulate visual dynamics and leverage them for downstream planning and control, rather than directly learning policies. [8750823] adopt a slightly different approach and propose a visual MPC controller that outputs a sequence of velocity commands based on the current image and a trajectory of subgoal images. Authors in [jacquet2024n] regress depth images and actions to collision probabilities which is then minimized within an MPC framework. Although theoretically, this approach can be extended to work with estimated depth, it is unclear how the associated noise will affect the efficiency of the overall pipeline. Moreover, we believe that predicting obstacle clearance distribution is easier than directly predicting collision probabilities, as the supervision data for the former can be easily obtained. Our design choice also allows us to use sophisticated statistical tools to capture collision risk from the clearance samples in an efficient manner.
@@
Improvement Over State-of-the-Art: Unlike end-to-end pipelines, our approach explicitly incorporates safety constraints by learning an action-conditioned collision model, placing us closer to [Bar_2025_CVPR], [hansen2024tdmpc2], [wu2022daydreamer], [8750823].However, in contrast to these cited works, we explicitly reason about uncertainty in collision prediction, and introduce a novel method to learn task-aware uncertainty. Moreover, we also learn optimal parameters of our risk metric to improve the efficacy of our approach. On the implementation side, most monocular navigation methods demonstrate results in simple environments like hallways. In contrast, we show, for the first time, robust and reliable navigation in cluttered settings typically requiring LiDAR or depth sensors.
@@ III-A Vision-Based Navigation as Risk-Aware Trajectory Optimization
$\mathbf{x}_{k+1}=f(\mathbf{x}_{k},\mathbf{u}_{k}),\ \mathbf{u}_{min}\leq\mathbf{u}_{k}\leq\mathbf{u}_{max}\forall k$ (1b)
where $c(\cdot)$ represents the state-dependent cost function. The vector $\mathbf{x}_{k}$ represents the state of the robot at time-step $k$. The vector $\mathbf{x}$ is the concatenation of the states at different $k$. The function $f$ represents the robot dynamics model. The control inputs are represented by $\mathbf{u}_{k}$. The vector $\mathbf{u}$ is formed by stacking $\mathbf{u}_{k}$ at different time-step $k$ respectively. The scalar function $d$ represents a learned probabilistic collision model, which takes as input the observations $\mathbf{o}$ and the control sequence $\mathbf{u}$ and outputs a distribution of worst-case obstacle clearance along the trajectory resulting from $\mathbf{u}$. We discuss the details of $d$ in the subsequent sections. The first term $c$ in (1a) minimizes the state cost, typically addressing path-following errors. The second term $r$ captures collision-risk based on the predictions of $d$. The last term $\left\|\mathbf{u}\right\|_{2}^{2}$ in (1a) penalizes large control inputs, and the weights $w_{i}$ tune the robot’s risk-seeking behavior. Control bounds are enforced through (1b). We construct a MPC feedback loop by solving (1a)-(1b) in a receding horizon manner from the current state.
@@ III-B Modeling Risk Through Chance-Constraints
$r(h(d(\mathbf{u},\mathbf{o})))=P(h(d(\mathbf{u},\mathbf{o}))\geq 0)$ (2)
Intuitively, our risk model captures the probability of collision constraints being violated for a given control sequence $\mathbf{u}$ and current observation $\mathbf{o}$. Moreover, the risk depends on the distribution characterized by the collision model $d$. Our aim is to first learn a suitable $d$ and then minimize the associated risk using (1a)-(1b).
A key challenge in using risk $r$ (2) is that the left-hand side does not have an analytical form if the predictive distribution of $d(\mathbf{u},\mathbf{o})$ departs significantly from Gaussian. Thus, in the next subsection, we present a tractable surrogate risk model which can produce the same effect as $r$.
@@ III-C Maximum Mean Discrepancy (MMD) as parameterized risk metric
In the deterministic scenario, driving $\overline{h}(d(\mathbf{u},\mathbf{o}))$ to zero will push $h(d(\mathbf{u},\mathbf{o}))$ to the feasible boundary. In the stochastic case, (3) maps $d(\mathbf{u},\mathbf{o})$ to a distribution of constraint residuals. Let $\overline{h}(d(\mathbf{u},\mathbf{o}))\sim p_{\overline{h}}$. Although we don’t know the parametric form for $p_{\overline{h}}$, we can be certain that its entire mass lies to the right of $\overline{h}=0$. Moreover, as $P(h(d(\mathbf{u},\mathbf{o}))\geq 0)$ approaches zero, $p_{\overline{h}}$ converges to a Dirac-Delta distribution $p_{\delta}$ [10979910]. In other words, one way of reducing risk is to minimize the difference between $p_{\overline{h}}$ and $p_{\delta}$. We formulate this distribution matching through the lens of Maximum Mean Discrepancy (MMD) [10.5555/2188385.2188410],[10979910] and use the following surrogate for the collision-risk proposed in [10979910], [11084873].
@@
The $K_{\lambda}$ is a positive definite function in RKHS known as the kernel function with the ”reproducing” property (6) [10.5555/2188385.2188410] , where $\left\langle\cdot\right\rangle$ represents the inner product. Throughout this paper, we have used Laplacian kernel for which $\lambda$ represents the kernel width.
@@ III-D Learning the Collision Model: The Baseline Approach
This section presents our approach for learning a probabilistic collision model $d(\mathbf{u},\mathbf{o})$ which quantifies the worst-case obstacle clearance along a given trajectory. We model it as a Gaussian random variable with mean $\mu_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$ and variance $\sigma_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$, both of which are designed as feed-forward neural networks parameterized by weight $\boldsymbol{\theta}$.Our design choice captures observation noise in $\mathbf{o}$, primarily arising from depth or point-cloud estimates of the vision foundation model [chua2018deep]. The detailed architecture is shown in Fig. 1. It consists of a pre-trained depth estimator that takes in RGB image and produces a depth-map. The depth-map is converted to a 2D point cloud and passed through PointNet++ [NIPS2017_d8bf84be] to get a feature embedding. We concatenate PointNet++ features with the initial state to get the full observation vector $\mathbf{o}$, which is then passed along with control sequence $\mathbf{u}$ to a Multi-Layer Perceptron (MLP) to get $\mu_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$ and $\sigma_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$. Both PointNet++ and MLP are trained in end-to-end manner using the following negative log-likelihood (NLL) function.
@@
1MM = Maximum number of iterations
@@ 28 $cost\leftarrow$ $c_{q}$, $\forall q$ in the ConstraintEliteSetConstraintEliteSet
35
36  0.1cm $({{}^{m+1}}\boldsymbol{\nu},{{}^{m+1}}\boldsymbol{\Sigma})\leftarrow$ Update distribution based on EliteSetEliteSet
@@
Algorithm 1 Sampling-Based Optimizer to Solve (1a)-(1b)
@@ III-E Learning the Collision Model: The Down-Stream Task-Aware Approach
The risk estimation head estimates collision constraint violations from $^{i}{d}$ ((8b)). Using these violations and the predicted kernel width $\lambda_{\boldsymbol{\theta}}$, the empirical MMD risk $r_{MMD}^{emp}$ is computed. This risk is then passed through an MLP to yield a 2D feature vector $\mathbf{z}$, which a softmax layer normalizes to $\hat{\mathbf{y}}$. The ground-truth label for $\hat{\mathbf{y}}$ is given by $\mathbf{y}$ and is computed through (8c)
@@
$^{i}h(^{i}d)=-{{}^{i}}{d}+d_{o},\ ^{i}\overline{h}(^{i}d)=\max(0,{{}^{i}}h(^{i}d))$ (8b)
@@
[…]
The architecture of Fig. 2 is trained with a combination of NLL and a cross-entropy loss (8d). The latter acts as an implicit task-aware regularization on the predicted $\sigma_{\boldsymbol{\theta}}(\mathbf{u},\mathbf{o})$.
@@ III-F Overall Approach
Alg. 1 presents our overall approach wherein we solve (1a)-(1b) using the trained neural worst-case obstacle clearance model as our collision predictor. We adopt a sampling-based optimizer based on [bhardwaj2022storm], wherein a batch of control sequences are drawn from a distribution which is gradually refined across iterations.
The algorithm initializes the sampling distribution (Line 2) and samples control inputs (Line 5), which are passed through a trained neural network (Line 6) to predict the mean $\mu_{\boldsymbol{\theta},q}$ and variance $\sigma_{\boldsymbol{\theta},q}$ associated with the distribution of obstacle clearances, as well as the kernel hyperparameter $\lambda_{\boldsymbol{\theta},q}$. These predictions generate $N$ distance samples $^{i}d_{q}$, constraint violations $^{i}\overline{h}_{q}$ (Line 7), and subsequently $r_{MMD,q}^{emp}$ (Line 8). State trajectories $\mathbf{x}_{q}(\mathbf{u}_{q})$ are then generated using the robot dynamics (Line 9). The $n_{c}$ lowest-risk samples form the $ConstraintElliteSet$ (Line 10), from which costs are evaluated and stored (Lines 11–13). Finally, the top $n_{e}$ samples form the $ElliteSet$ (Line 14) used to update the sampling distribution (Line 15) following [bhardwaj2022storm].